OpenStudy (anonymous):

Stoila needs to determine the wquation of a line passing through the point (-6, -2) that is perpendicular to y=2x-4. What is the equation?

OpenStudy (btaylor):

the slope would be the negative reciprocal to be perpendicular, so it is now -0.5 then use the equation: \[y-y_{1} = m(x-x_{1})\] Plugging in the values: \[y+2=-0.5(x+6)=-.5x-3\] Solve for y, to get \[y = -.5x-5\]

OpenStudy (anonymous):

working it out now....thank you f your help!